### v2

Abstract
In previous articles we argued that if both influence at a distance and super-determinism are considered unacceptable to explain the results of loophole-free Bell tests, a third way is available, called “predictive incompleteness”: it tells that the usual quantum state $\psi$ is incomplete, as long as the measurement context has not been specified. This is in agreement with the general quantum framework called CSM (Contexts, Systems and Modalities), which provides a unified approach for considering quantum systems within classical contexts. In this article we further illustrate how predictive incompleteness is working, by looking at other ideas based on either the Kolmogorovian Censorship (KC), or on superdeterminism and supermeasurements. We show that these possibilities are very advantageously replaced by the use of predictive incompleteness.
I Introduction.
In the scientific literature there are still debates to decide whether quantum mechanics (QM) is contextual ContextualReview ; ContextualPG or noncontextual histories , and whether quantum probabilities fit within the usual Kolmogorovian framework LZ1 ; LZ2 or not holik ; svozil ; book . To some extent, these debates can be considered as matters of definitions, so it is useful to spell out the explicit or implicit assumptions that are behind these different views. A milestone in this debate is the Kolmogorovian Censorship (KC), stating that quantum probabilities can be identified with classical, Kolmogorovian probabilities when the measurement context has been specified, or more generally when a classical probability distribution of measurement contexts has been specified. The KC sounds rather obvious on a physical intuitive basis, and technically it has been demonstrated for a countable number of measurement contexts proof ; redei . It also has some relationship with superdeterminism or supermeasurements of Refs. Sabine ; Palmer as it will appear more clearly below.
II Kolmogorovian Censorship and Bell experiments.
Looking in more details, there is some controversy on the physical meaning of the KC, as we will explain now. In particular, Szabó et al LZ1 ; LZ2 claim that QM probabilities are Kolmogorovian, not only in a single context, but also in a loophole-free Bell experiment, where a random choice among four different contexts is implemented aa . Then they conclude that Bell’s inequalities are either irrelevant (when considering the four contexts separately), or not violated, when considering the four contexts together: since each one has a 1/4 probability to occur, Bell’s $S$ parameter (see Annex) is reduced from $2\sqrt{2}$ down to 2/$\sqrt{2}/2$, that is below 2 as it would be expected classically.
More precisely, it is uncontroversial that the probabilities predicted by QM in a given context are Kolmogorovian. and the problem arises when trying to “put together” probabilities predicted by QM in different contexts. Said otherwise, in any given context it is possible to build a Kolmogorovian probability distribution, or equivalently a hidden variable theory; however this distribution must be adapted to the chosen context.
In a loophole-free Bell test aa , the remote random choices of measurements (polarizers orientations) are designed to forbid that this matching may be done by a relativistically causal transmission between the source and the detectors. Despite these barriers, the matching might be obtained either by an instantaneous influence between the source and the detectors (explicit nonlocality), or by assuming that it is pre-established before the actual experiment takes place (super-determinism). One has then the following possible options:
  1. 1.
    Szabó et al LZ1 ; LZ2 claim that the only relevant way to interpret QM probabilities in different contexts is to consider a classical probability distribution over these different contexts. This corresponds by construction to a feasible experiment, typically a loophole-free Bell experiment, including the random choice among four different contexts aa . Then the global probability distribution is Kolmogorovian, and Bell’s inequalities (BI) are not violated, because as written above the probabilities in each context are divided by four, as well as the resulting $S$ value. However a drawback of this approach is that the pre-established matching between the source and the measurements is still required, through some kind of “global determinism” durt . Therefore, though BI are not explicitly violated, the basic physical problem of the origin of the matching is still present.
  2. 2.
    Rather than considering a probability distribution over the contexts, one may calculate the correlation function in each context, which is the very idea of a loophole-free Bell test, and gather them in Bell’s $S$ value. However, one may argue that the four correlations functions correspond to four different incompatible experiments, and thus bringing their results together is counterfactual, because they cannot be measured simultaneously; then BI cannot be demonstrated. This is a standard answer to the problem, but without further explanation it sounds more dogmatic than scientific; see III.B for an attempted improvement.
  3. 3.
    Still an alternative way is to admit that the results in the four contexts can be combined, as it would be the case classically since they apply to the same system; then one gets Bell’s inequalities, which are experimentally violated, so one should explain why. We consider that this is a meaningful question, and there are basically three options, spelled out in detail in entropy :
    3a - in the spirit of option 1 above, the way to violate Bell’s inequalities is by admitting that the system’s parameters and the orientations of the polarizers are not independent variables, despite the fact that these orientations are chosen randomly and independently at a large distance. This can be obtained either by admitting superdeterminism (i.e. denying the possibility of independent random choices durt ), or by admitting a non-local influence between the source and the measurements entropy . These two options have recently been shown to be equivalent blasiak , and they are in our opinion equally undesirable - though they are matter of ontological choice, and cannot be proven wrong.
    3b - the third option, known as predictive incompleteness entropy ; jarrett , is to recognize that the quantum state by itself is not enough to specify the measured probability distribution, as long as the context has not be specified. This contradicts predictive completeness (also called outcome independence), which is a required hypothesis for Bell’s theorem, and therefore BI cannot be demonstrated. This conclusion corresponds to the detailed analysis presented in entropy , and also the more general framework presented in completing to “complete” the usual quantum state by specifying the measurement context. These two papers, as well as the arguments above, are consistent within the general quantum framework called CSM (Contexts, Systems and Modalities) CSM1 ; CSM2 .
III Discussion.
III.1 Kolmogorovian censorship vs predictive incompleteness.
From the above it should be clear that the ontologies underlying either Szabó’s position or CSM are quite different. Szabó et al claim that QM probabilities are Kolmogorovian and that QM can ultimately be seen as a (super)deterministic theory. On the other hand, CSM stipulates that QM is fundamentally non-deterministic, due to the conjunction of quantization and contextuality Gleason ; Uhlhorn , and that quantum probabilities are non-Kolmogorovian, unless restricted to a single context according to the KC.
This is why the (non-Kolmogorovian) predictive incompleteness of $\psi$ is useful: it leaves enough freedom so that the independent choice of the measurement can contribute to the determination of the observed probability distributions, avoiding both nonlocality and superdeterminism. It is worth emphasizing again, as spelled out in entropy , that predictive incompleteness makes no sense in classical physics, and appears as a specific quantum feature - thereby able to consistently explain the violation of Bell’s inequalities.
III.2 What about non-measurable subsets ?
Let’s note that the main proof in LZ2  is based on a theorem in Pitowsky’s book Pbook , which speaks also about the violation of Bell’s inequalities using non-measurable subsets of local hidden variables (Chapter 5 in Pbook ). This kind of ideas were developed more recently Sabine ; Palmer , and are very controversial Scott ; Matteus . We don’t want to enter in this debate here, but we note the following sentence by Palmer Palmer : “Importantly, this means a non-conspiratorial interpretation of (our approach) implies that physical theory does not have the post-hoc property of counterfactual definiteness. (…) This important point seems to have been lost in more recent discussions of Bell’s Theorem.”
This gives a major hint about the meaning of Palmer’s approach: if counterfactual definiteness - that is, treating in statistical calculations the results of unperformed measurements on an equal footing as actual results - is rejected, then Bell’s inequalities don’t hold any more (option 2. above); this is well known and has not ‘been lost in more recent discussions of Bell’s Theorem’. Then the work in Sabine ; Palmer may be seen as a possible way to justify a rejection of counterfactual definiteness Hance , but there is a much simpler one : counterfactual definiteness is not compatible with predictive incompleteness, see Annex and entropy . Actually Ref. Sabine gives a good definition of predictive incompleteness: it is “the lack of information about the measurement outcomes in the wave-function, combined with the fact that an observation in one location can tell us something about the measurement outcome in another location”. Then Sabine is trying to explain this by using supermeasurements, but here we get it without using any non-measurable set of hidden variables - as a consequence of the CSM contextual quantization postulates CSM1 ; CSM2 ; Gleason ; Uhlhorn . Again, acknowledging that the quantum state vector is predictively incomplete may appear disturbing, but it makes perfect sense and removes many “paradoxes” in the CSM framework.
IV Conclusions.
Again, the KC in a single context is easily integrated in the CSM framework, but one should not conclude that quantum probabilities are Kolmogorovian in a classical sense: this would be true in classical physics, because there is only one universal context, but this fails in quantum physics, because there is a continuous infinity of different, incompatible contexts. This point of view is implicit in textbook quantum mechanics, and it is made explicit in the CSM framework, by using operator algebra and infinite tensor products. This framework allows a contextual unification of classical and quantum physics, within a unique macroscopic physical world MP1 ; MP2 ; MP3 ; enigma .
One may notice that Rédei redei (see also gomori ) considers that the KC is problematic, in particular because it implies that “probabilities are thus not features of quantum systems in and of themselves, they are features that only manifest themselves upon measurement. Philosophers (or physicists) with a robust realist conviction may find unattractive this strongly instrumentalist flavor of interpretation of quantum probability forced upon us by the KC.” In the CSM approach, quantum probabilities also get a meaning only upon measurement, due to the predictive incompleteness of $\psi$, and they are also genuinely non-classical probabilities. Nevertheless, CSM is based on physical realism, as the statement that the purpose of physics is to study entities of the natural world, existing independently from any particular observer’s perception, and obeying universal and intelligible rules completing ; CSM1 ; CSM2 . We claim therefore that even with a “robust realist conviction” one can accept that $\psi$ is predictively incomplete - as the best way to make sense of this conundrum, within the framework of contextual objectivity CO2002 .
As a final remark, a significant challenge in quantum reconstructions is to state clearly what are the hypotheses or postulates, and what are their consequences. It should also been made clear whether one looks for a fully deductive reasoning, or a partially inductive one, that is an “Inference to the Best Explanation” Gleason ; Uhlhorn . In the CSM point of view a fully deductive approach does not fit, since, quoting Landau landau , ‘quantum mechanics (…) contains classical mechanics as a limiting case, yet at the same time it requires this limiting case for its own formulation’. Then what is desired is not a deduction of the Laws of Nature from some postulates in a mathematical sense, but rather a fully consistent construction, including both classical and quantum physics from the beginning, and clearly separating experimentally based evidence from its mathematical description MP1 ; MP2 ; MP3 ; enigma .
Acknowledgments. The author thanks Mathias Van Den Bossche and Olivier Ezratty for relevant comments on this article, and for many interesting discussions.
[…]
On the other hand, if the normalization of the correlation coefficients is done by summing all counts for all possible orientations and results, one may consider that $x$ and $y$ are also issued from independent random processes in variable spaces $X$ and $Y$, as it is done in loophole-free Bell tests aa . Then the global probability writes
$P(xa,yb)=\sum_{\lambda\in\Lambda,x\in X,y\in Y}P(ab|xy\lambda)P(\lambda)P(x)P(y)$ (1)
where $P(ab|xy\lambda)$ is given by standard formulas entropy . Taking $\Lambda=\{\lambda\}$, $X=\{x,x^{\prime}\}$ with $P(x)=P(x^{\prime})=1/2$, $Y=\{y,y^{\prime}\}$ with $P(y)=P(y^{\prime})=1/2$ as in a usual Bell test, one gets $P(xa,yb)=P(ab|xy\lambda)/4$.
Correspondingly, the random variable $(xa,yb)$ may take 16 mutually exclusive values, not 4, and Bell’s inequalities cannot be written anymore. This is because Bell’s reasoning requires to calculate the correlation functions $E(x,y)=\langle ab\rangle_{x,y}$ by using $P(ab|xy)$, not $P(xa,yb)$, so that the four different measurements apply to the same sample space $\Lambda$. This means implicitly that $\lambda$ completely carries the pairs’ properties, and the measurement results can be predicted from the knowledge of $\lambda$ alone, as it would be the case in classical physics. But this is counterfactual with respect to the quantum approach, where $\{\lambda,x,y\}$, $\{\lambda,x,y^{\prime}\}$, $\{\lambda,x^{\prime},y\}$, $\{\lambda,x^{\prime},y^{\prime}\}$ are four different situations which should not be merged within a single $S$ value, contrary to Bell’s reasoning Bell ; BCH . This is another way to tell that $\psi$ is predictively incomplete, and requires a context specification to be turned into an actual probability distribution.